# Prevention is better than cure? Feedback from high specific energy winds in cosmological simulations with Arkenstone

### v1

@@ 2.6 Mass and energy loadings
Figure 1: Mass (top panel) and energy (bottom panel) loadings for all Arkenstone simulations in this paper, in comparison to scalings from TNG (from the $z=0$ fits to $\eta_{\rm M}$ and $v_{\rm w}$ in Pillepich et al., 2018a). Both mass and energy loadings in Arkenstone runs are generally much lower than in TNG. Throughout most plots in this paper, shades of orange show simulations with constant $\eta_{\mathrm{E}}$ values and fixed $\eta_{\mathrm{M}}=0.3$; pink/purple lines show simulations with constant $\eta_{\mathrm{M}}$ values and fixed $\eta_{\mathrm{E}}=0.3$; shades of blue show simulations with variable energy loading with a fixed $\eta_{\mathrm{M}}=0.3$. We note that the input loading values actually applied in the simulations for the latter and TNG have an associated scatter, due to the use of the locally measured DM velocity dispersion (and metallicity in the case of TNG) in their calculation.
In this paper we present results from a total of 10 cosmological simulations. In all results we compare to the output of the existing TNG model (but without MHD), ran on the same simulation box. Specifically, we have three runs with a fixed mass loading $\eta_{\mathrm{M}}=0.3$, and fixed energy loadings of $\eta_{\mathrm{E}}=0.1$, 0.3, and 0.9 (shown in all plots in shades of orange). This is complemented with three runs with a fixed energy loading $\eta_{\mathrm{E}}=0.3$, and fixed mass loadings of $\eta_{\mathrm{M}}=0.1$, 0.3, and 0.9 (shown in pink/purple). We note that the run with $\eta_{\mathrm{E}}=0.3$ and $\eta_{\mathrm{M}}=0.9$ catastrophically failed to regulate star formation after $z\sim 0.3$ (see Section 3.5), and became too computationally expensive to reach $z=0$. This run is therefore not displayed in figures only showing data from $z=0$.
@@
The slope $\alpha$ is the expected scaling we wish to realise with halo mass; the additional factor of $3$ comes due to the relationship $\sigma_{\rm DM,1D}\propto M_{\rm h}^{1/3}$. $\sigma_{\rm DM,1D,0}$ is the characteristic DM velocity dispersion for a $10^{11}$ M halo, which we take to be $51$ km s-1. At this halo mass we then take the expected energy loading for a $10^{11}$ M halo to be $\eta_{\mathrm{E,norm}}=0.4309$ (equivalent to a normalisation of 0.2 for a halo mass of $10^{12}$ M when $\alpha=-1/3$). Equation (14) is the input energy loading to simulation these variable energy loading simulations, which then approximately follows the halo mass scaling
$\eta_{\mathrm{E}}=\eta_{\mathrm{E,norm}}\left(\frac{M_{\rm h}}{10^{11}\,\mathrm{M}_{\odot}}\right)^{\alpha}.$ (15)
Our three runs forming the final group still include a fixed mass loading $\eta_{\mathrm{M}}=0.3$, but include this parametrisation of a variable energy loading with slopes $\alpha=-1/6,-1/3,$ and $-1/2$. We cap the energy loading in these simulations at $\eta_{\mathrm{E,max}}=1$. All runs in this paper and their basic properties are summarised in Table 1.
We show our loadings for both mass (top panel) and energy (bottom panel) in Fig. 1, compared to the input values for TNG (from the fits of Pillepich et al., 2018a). In TNG, the energy loading (dependent on metallicity) and launch velocity (dependent on redshift and with a minimum value of $350\,{\rm km\,s^{-1}}$) are model inputs, and this then sets the mass loading. The values for our variable energy loading simulations and for TNG actually show a scatter around the analytic lines shown in Fig. 1, induced by scatter in the DM velocity dispersion. It is immediately obvious in the top panel that the mass loadings of all Arkenstone runs are dramatically smaller than for TNG in low mass galaxies – the central mass loading value used in most Arkenstone runs is 5, 34, and even 100 times lower than TNG for haloes of $10^{12}$, $10^{11}$, and $10^{10}$ M, respectively.
The bottom panel indicates how energy loadings are also high in TNG, due to the requirement of launching large amounts of mass out of the galaxy. The lowest energy loading in TNG, reached at high metallicities, is 0.9 – the highest fixed value we study here. Even our variable runs, which are capped at $\eta_{\mathrm{E}}=1$, still have energy loadings at least 3.6 times lower than TNG for low mass galaxies.
@@ 3 Results
3.1 First look
@@
The differences between the simulations are emphasised in Fig. 3, which shows the ratio of the two maps in Fig. 2. Here, blue regions highlight where the gas surface density is lower in Arkenstone than TNG, which occurs close to many galaxies in the simulation volume. Orange regions show where gas surface densities are higher in the Arkenstone run, which generally happens on the outskirts of haloes and into the IGM. We note that many of the changes seen in gas density here materialise well outside the virial radii of haloes, affecting the outer CGM and IGM, showing how high specific energy winds can have important impacts far from the galaxies that launch them. Detailed properties of the IGM in Arkenstone simulations will be studied in future work, although we note that a generic prediction is a higher IGM gas fraction, potentially in agreement with recent observations (Connor et al., 2024).
@@ 3.2 Galaxy properties at $z=0$
Figure 4: Stellar mass functions at $z=0$ for simulations with: 1) constant energy loadings and fixed mass loading of $\eta_{\rm M}=0.3$ (top panel), 2) constant mass loadings and fixed energy loading of $\eta_{\rm E}=0.3$ (middle panel), and 3) simulations with variable energy loading with different slopes $\alpha$ and a fixed mass loading of $\eta_{\rm M}=0.3$ (bottom panel). Each panel also shows the result from TNG (black solid line) and observational data (black dotted line) from Shuntov et al. (2022). Lines shown are the median of the data, smoothed over the nearest two bins. For the hot, light winds modelled by Arkenstone, energy loading affects the normalisation and slope of the SMF much more than mass loading.
Figure 5: Ratios of stellar mass to halo mass at $z=0$ as a function of halo mass, normalised by the universal baryon fraction ($f_{\mathrm{B}}=0.16$, Planck Collaboration et al., 2016). Median lines have the same styles as Fig. 4, with observational data also from Shuntov et al. (2022). For the largest hales in our box we plot individual scatter points, and change our median lines to be dashed. This is to emphasise that results at high masses should be treated with caution, due to the lack of massive haloes and the dominance of the TNG BH model here. Fixed energy loadings tend to overproduce stars at the low mass end, which can be alleviated by introducing a variable energy loading.
From this Section to Section 3.5, we show a number of population-level results for the properties of galaxies and their gaseous haloes in our simulated boxes. For all plots except the SFR density, solid lines show the median of each quantity, smoothed over the nearest two points for clarity.
In all panels, the solid black line shows the quantity for the TNG model run on the same box. For Arkenstone runs, lines with shades of orange show the results for runs with a fixed mass loading ($\eta_{\mathrm{M}}=0.3$) and several constant energy loading values. Shown in pink and purple are runs with a fixed energy loading ($\eta_{\mathrm{E}}=0.3$) and several constant mass loading values. Finally, lines in shades of blue show results for runs with a variable energy loading. Dotted black lines in each panel show observational constraints where available, though we note that a more detailed comparison of full mock observations is deferred to future work.
We further note that due to our limited box size of (36.936.936.9\,36.9Mpc)3, our results at high stellar and halo masses are limited by a lack of large objects. For this reason, and the uncertain interplay between Arkenstone and the TNG BH model, we suggest caution in interpreting results at these masses. We show individual data points for the largest few haloes in the box (above $M_{\rm h}\sim 10^{13}\,$), and plot the medians as dashed lines.
@@ 3.2.1 Stellar Mass Functions
Firstly in Fig. 4, we show the SMFs at $z=0$ for all of our runs. The solid black line in all panels shows the SMF for the TNG model, the parameters of which were tuned to reproduce this quantity. In the top panel, showing effects of changing energy loading while holding mass loading constant, the cumulative feedback effect of energy loading of a hot wind is very clear. The palest line shows the lowest energy loading, $\eta_{\mathrm{E}}=0.1$, which is evidently ineffective in regulating star formation across a wide range of galaxy masses. The intermediate energy loading of $\eta_{\mathrm{E}}=0.3$ is shown in the middle orange colour, which brings the SMF down closer to the observations, especially for galaxies at a stellar mass of $\sim 10^{10}$ M. This run still shows an excess of galaxies with $M_{\star}<10^{9}$ M, suggesting a fixed energy loading of 0.3 is not enough to regulate these dwarf galaxies. We discuss this further below. The darkest line in the top panel shows the highest energy loading, $\eta_{\mathrm{E}}=0.9$, which overregulates star formation in galaxies with $M_{\star}>10^{9}$ M, but produces a distribution closer to the observations for $M_{\star}<10^{9}$ M. This points to the need for a higher energy loading in lower mass systems.
The middle panel of Fig. 4 shows simulations with changes in mass loading for a fixed energy loading of $\eta_{\mathrm{E}}=0.3$. The effect of these changes in mass loading is much smaller than in energy loading, though the amplitude of the SMF does decrease slightly with decreasing mass loading (i.e. less mass ejection results in less star formation). As we discuss later, this is due to a lower mass loading leading to a hotter, faster wind, resulting in a higher level of preventative feedback due to CGM heating. This is the opposite behaviour to regulation via ejective feedback, where increasing the mass loading tends to lower the star formation by reducing the amount of material in the ISM (as long as there is still enough energy to throw the mass far enough). As a reminder, the run with $\eta_{\rm M}=0.9$ (and $\eta_{E}=0.3$) completely failed to regulate star formation at low redshifts and did not reach $z=0$ – this wind has too low a specific energy to provide effective preventative feedback while also having too small a mass loading to act as an efficient source of ejective feedback.
@@
3.2.2 Stellar Mass Halo Mass Relation
@@
A similar account emerges in the SMHM relation in Fig. 5, here normalised by the cosmic baryon fraction. Recall that this relation was used to calibrate the TNG model. The different (constant) values of energy loading in the top panel show the strong sensitivity of the SMHM to the value of $\eta_{\rm E}$, with the gradient getting steeper with higher energy loading. The lowest energy loading clearly leads to a significant overproduction of stars at the low-mass end, whereas the high-mass end closely follows TNG, likely due to the dominance of the AGN feedback model. An energy loading of 0.3 reduces the stellar mass across all halo masses, most notably below $10^{12}$ M, though there is still an overproduction of stars in the dwarf regime. The highest energy loading of 0.9 dramatically and too effectively reduces the stellar mass of many galaxies – the SMHM relation lies nearly an order of magnitude below the observational data at a Milky Way halo mass of $\sim\!10^{12}$ M. Though in the dwarf regime the SMF is closer to the observations, again pointing towards needing higher energy loading in low mass galaxies. Notably, the higher energy loadings also lead to somewhat reduced stellar masses in the largest haloes in our box, though this could be due to the effect of stellar feedback on the growth of BHs and their subsequent feedback.
@@
The mass loading of high specific energy winds can also affect the number of stars formed in lower mass haloes, but much less significantly than energy loading (we note that the gradient of this relation is strongly dependent on energy loading). Lower mass loadings, via hotter, faster winds, lead to less stars forming in lower mass haloes. We note that this ceases to be the case at halo masses above $\sim\!10^{12.5}\,{\rm M}_{\odot}$, though as we have discussed, interpretation here is complicated by the AGN feedback.
@@
As with the SMF, having a variable energy loading with halo mass brings the SMHM for Arkenstone into better agreement with observations, by further decreasing the stellar masses of low mass haloes. The slope of the variable energy loading directly affects the slope of the SMHM at low masses, with the steepest relation leading to more suppression below $\sim\!10^{11}$ M, and less above that. Our results seem to favour a steeper gradient of $\eta_{\rm E}$, closer to that of Carr et al. (2023) than that of Voit et al. (2024b). At the high mass end, the energy loading becomes small, and so the SMHM relation is set entirely by the TNG BH model.
3.2.3 Stellar Mass Black Hole Mass Relation

### v2

@@ 2.6 Mass and energy loadings
Figure 1: Mass (top panel) and energy (bottom panel) loadings at $z=0$ for all Arkenstone simulations in this paper, in comparison to scalings from TNG (from the $z=0$ fits to $\eta_{\rm M}$ and $v_{\rm w}$ in Pillepich et al., 2018a). Both mass and energy loadings in Arkenstone runs are generally much lower than in TNG. Throughout most plots in this paper, shades of orange show simulations with constant $\eta_{\mathrm{E}}$ values and fixed $\eta_{\mathrm{M}}=0.3$; pink/purple lines show simulations with constant $\eta_{\mathrm{M}}$ values and fixed $\eta_{\mathrm{E}}=0.3$; shades of blue show simulations with variable energy loading with a fixed $\eta_{\mathrm{M}}=0.3$. We note that the input loading values actually applied in the simulations for the latter and TNG have an associated scatter, due to the use of the locally measured DM velocity dispersion (and metallicity in the case of TNG) in their calculation. Arkenstone runs with variable energy loadings are capped at an $\eta_{\mathrm{E}}=1$.
In this paper we present results from a total of 10 cosmological simulations. We compare the predictions from the Arkenstone model to the output of the existing TNG model (but without MHD), which we ran using the same initial conditions and simulation configuration. Specifically, we have three runs with a fixed mass loading $\eta_{\mathrm{M}}=0.3$, and fixed energy loadings of $\eta_{\mathrm{E}}=0.1$, 0.3, and 0.9 (shown in all plots in shades of orange). This is complemented with three runs with a fixed energy loading $\eta_{\mathrm{E}}=0.3$, and fixed mass loadings of $\eta_{\mathrm{M}}=0.1$, 0.3, and 0.9 (shown in pink/purple).
@@
The slope $\alpha$ is the expected scaling we wish to realise with halo mass; the additional factor of $3$ comes due to the relationship $\sigma_{\rm DM,1D}\propto M_{\rm h}^{1/3}$ (at fixed redshift). $\sigma_{\rm DM,1D,0}$ is the characteristic DM velocity dispersion for a $10^{11}$ M halo at $z=0$, which we take to be $51$ km s-1. At this halo mass we then take the expected energy loading to be $\eta_{\mathrm{E,norm}}=0.4309$ (equivalent to a normalisation of 0.2 for a halo mass of $10^{12}$ M at $z=0$ when $\alpha=-1/3$). Equation (14) is the input energy loading for these variable energy loading simulations, which then approximately follows the halo mass scaling
$\eta_{\mathrm{E}}\approx\eta_{\mathrm{E,norm}}\left[\frac{M_{\rm h}}{10^{11}\,\mathrm{M}_{\odot}}\left(\frac{H(z)}{H_{0}}\right)^{1/3}\right]^{\alpha}.$ (15)
Our three runs forming the final group still include a fixed mass loading $\eta_{\mathrm{M}}=0.3$, but include this parametrisation of a variable energy loading with slopes $\alpha=-1/6,-1/3,$ and 1/2-1/2. For simplicity, we generally drop the redshift dependence in Eq. 15 when discussing runs with this scaling. All runs in this paper and their basic properties are summarised in Table 1.
We additionally cap the energy loading in all Arkenstone simulations at $\eta_{\mathrm{E,max}}=1$. Realistically, the actual amount of energy available to drive winds out of the galaxy will be less than $\eta_{\mathrm{E,max}}$, as work will need to be done both to inflate bubbles and drive turbulence in the ISM. An energy loading $\eta_{\mathrm{E}}>1$ is clearly too high, which is why we cap our runs at $\eta_{\mathrm{E,max}}$, but to better understand the true amount of energy available to drive winds we must turn to high resolution, more idealised simulations (e.g. TIGRESS, Kim & Ostriker, 2018), loadings from which we will investigate in future work.
We show our loadings for both mass (top panel) and energy (bottom panel) in Fig. 1, compared to the input values for TNG (from the fits of Pillepich et al., 2018a). In TNG, the energy loading (dependent on metallicity) and launch velocity (dependent on redshift and with a minimum value of $350\,{\rm km\,s^{-1}}$) are model inputs, and this then sets the mass loading. The values for our variable energy loading simulations and for TNG actually show a scatter around the analytic lines shown in Fig. 1, induced by scatter in the DM velocity dispersion (and due to the metallicity dependence, in the case of TNG). It is immediately obvious from the top panel that the mass loadings of all Arkenstone runs are dramatically smaller than for TNG in low mass galaxies – the central mass loading value used in most Arkenstone runs is 5, 34, and even 100 times lower than TNG for haloes of $10^{12}$, $10^{11}$, and $10^{10}$ M, respectively.
The bottom panel of Fig. 1 shows that energy loadings are also high in TNG, as a direct consequence of the high mass loadings and fixed thermal energy fraction of the winds. Even the maximum energy loading in our variable runs, which are capped at $\eta_{\mathrm{E}}=1$, still have energy loadings at least 3.6 times lower than TNG for low mass galaxies.
@@ 3 Results
In this section we present our simulations results from all cosmological simulations, and as a comparison we show results from Arkenstone alongside those from the existing TNG model (which is calibrated to match certain observations). Where available, we also show a comparison with physical quantities estimated from observational data for selected quantities. We note, however, that in this work we do not forward model our simulation predictions into the observational plane, so the comparisons with observational data are intended to be qualitative and should be treated with caution.
3.1 Large-scale gas distribution
@@
The differences between the simulations are emphasised in Fig. 3, which shows the ratio of the two maps in Fig. 2. Here, blue regions highlight where the gas surface density is lower in Arkenstone than TNG, which occurs close to many galaxies in the simulation volume. Orange regions show where gas surface densities are higher in the Arkenstone run, which generally happens on the outskirts of haloes and into the IGM. We note that many of the changes seen in gas density here materialise well outside the virial radii of haloes, affecting the outer CGM and IGM, showing how high specific energy winds can have important impacts far from the galaxies that launch them. Detailed properties of the IGM in Arkenstone simulations will be studied in future work, although we note that a generic prediction is a higher IGM gas fraction, potentially in agreement with recent observations (Connor et al., 2025).
@@ 3.2 Galaxy properties at $z=0$
Figure 4: Stellar mass functions at $z=0$ for simulations with: 1) constant energy loadings and fixed mass loading of $\eta_{\rm M}=0.3$ (top panel), 2) constant mass loadings and fixed energy loading of $\eta_{\rm E}=0.3$ (middle panel), and 3) simulations with variable energy loading with different slopes $\alpha$ and a fixed mass loading of $\eta_{\rm M}=0.3$ (bottom panel - we remind the reader that we have dropped the redshift dependence on $\eta_{\rm E}$ in this panel for convenience). Each panel also shows the result from TNG (black solid line) and observational data from Bernardi et al. (2017) (black dotted line) and Driver et al. (2022) (black dashed line). Lines shown are the median of the data, smoothed over the nearest two bins. For the hot, light winds modelled by Arkenstone, energy loading affects the normalisation and slope of the SMF much more than mass loading.
Figure 5: Ratios of stellar mass to halo mass at $z=0$ as a function of halo mass, normalised by the universal baryon fraction ($f_{\mathrm{B}}=0.16$, Planck Collaboration et al., 2016). Median lines have the same styles as Fig. 4, with comparison data from Behroozi et al. (2019). For the largest haloes in our box we plot individual scatter points, and change our median lines to be dashed. This is to emphasise that results at high masses should be treated with caution, due to the lack of massive haloes and the dominance of the TNG BH model here. Fixed energy loadings tend to overproduce stars at the low mass end, which can be alleviated by introducing a variable energy loading.
From this Section to Section 3.5, we show a number of population-level results for the properties of galaxies and their gaseous haloes in our simulated boxes. For all plots except the SFR density and stellar mass function, solid lines show the median of each quantity, smoothed over the nearest two points for clarity.
In all panels, the solid black line shows the quantity for the TNG model run with the same initial conditions and with the same simulation configuration. For Arkenstone runs, lines with shades of orange show the results for runs with a fixed mass loading ($\eta_{\mathrm{M}}=0.3$) and several constant energy loading values. Shown in pink and purple are runs with a fixed energy loading ($\eta_{\mathrm{E}}=0.3$) and several constant mass loading values. Finally, lines in shades of blue show results for runs with a variable energy loading. Other lines in each panel show observational constraints where available, though we note that a more detailed comparison with observations is deferred to future work.
We further note that due to our limited box size of (36.936.9\,Mpc)3, our results at high stellar and halo masses are limited by poor statistics and large cosmic variance for large objects. For this reason, and the uncertain interplay between Arkenstone and the TNG BH model, we suggest caution in interpreting results at these masses. We show individual data points for the largest few haloes in the box (above $M_{\rm h}\sim 10^{13}\,$M), and plot the medians as dashed lines.
@@ 3.2.1 Stellar mass functions
In Fig. 4, we show the SMFs at $z=0$ for all of our runs. The solid black line in all panels shows the SMF for the TNG model, the parameters of which were tuned to reproduce this quantity (although using different observational data). In the top panel, showing effects of changing energy loading while holding mass loading constant, the cumulative feedback effect of energy loading of a hot wind is very clear. The palest line shows the lowest energy loading, $\eta_{\mathrm{E}}=0.1$, which is evidently ineffective at regulating star formation across a wide range of galaxy masses. The intermediate energy loading of $\eta_{\mathrm{E}}=0.3$ is shown in the middle orange colour, which brings the SMF down closer to the observations, especially for galaxies at a stellar mass of $\sim 10^{10}$ M. This run still shows an excess of galaxies with $M_{\star}<10^{9}$ M, suggesting a fixed energy loading of 0.3 is not enough to regulate these dwarf galaxies. We discuss this further below. The darkest line in the top panel shows the highest energy loading, $\eta_{\mathrm{E}}=0.9$, which overregulates star formation in galaxies with $M_{\star}>10^{9}$ M, but produces a distribution closer to the observations for $M_{\star}<10^{9}$ M. This points to the need for a higher energy loading in lower mass systems.
The middle panel of Fig. 4 shows simulations with changes in mass loading for a fixed energy loading of $\eta_{\mathrm{E}}=0.3$. The effect of these changes in mass loading is smaller than the changes from varying the energy loading, though the amplitude of the SMF does decrease slightly with decreasing mass loading (i.e. less mass ejection results in less star formation). As we discuss later, this is due to a lower mass loading leading to a hotter, faster wind, resulting in a higher level of preventative feedback due to CGM heating. This is the opposite behaviour to regulation via ejective feedback, where increasing the mass loading tends to lower the star formation by reducing the amount of material in the ISM (as long as there is still enough energy to throw the mass far enough). The run with $\eta_{\rm M}=0.9$ (and $\eta_{E}=0.3$) failed to regulate star formation at low redshifts across a wide halo mass range – this wind has too low a specific energy to provide effective preventative feedback while also having too small a mass loading to act as an efficient source of ejective feedback.
@@
3.2.2 Stellar mass – halo mass relation
@@
A similar picture emerges in the SMHM relation in Fig. 5, here normalised by the cosmic baryon fraction. Recall that this relation (although different data) was used to calibrate the TNG model. The different (constant) values of energy loading in the top panel show the strong sensitivity of the SMHM to the value of $\eta_{\rm E}$, with the gradient getting steeper with higher energy loading. The lowest energy loading clearly leads to a significant overproduction of stars at the low-mass end, whereas the high-mass end closely follows TNG, likely due to the dominance of the AGN feedback model. An energy loading of 0.3 reduces the stellar mass across all halo masses, most notably below $10^{12}$ M, though there is still an overproduction of stars in the dwarf regime. The highest energy loading of 0.9 dramatically and too effectively reduces the stellar mass of many galaxies – the SMHM relation lies nearly an order of magnitude below the observational data at a Milky Way halo mass of $\sim\!10^{12}$ M. In the dwarf regime the SMF is closer to the observations, again pointing towards the need for higher energy loading in low mass galaxies. Notably, the higher energy loadings also lead to somewhat reduced stellar masses in the largest haloes in our box - this is most likely due to strong winds affecting the progenitors of these objects at early times.
@@
The mass loading of high specific energy winds can also affect the number of stars formed in lower mass haloes, but much less significantly than energy loading (we note that the gradient of this relation is strongly dependent on energy loading). Lower mass loadings, via hotter, faster winds, lead to fewer stars forming in lower mass haloes. We note that this difference gets smaller at halo masses above $\sim\!10^{12.5}\,{\rm M}_{\odot}$, though as we have discussed, the interpretation here is complicated by the AGN feedback.
@@
As with the SMF, adopting a variable energy loading with halo mass brings the SMHM for Arkenstone into better agreement with observations, by further decreasing the stellar masses of low mass haloes. The slope of the variable energy loading directly affects the slope of the SMHM at low masses, with the steepest relation leading to more suppression below $\sim\!10^{11}$ M, and less above that. Our results seem to favour a steeper gradient of $\eta_{\rm E}$, closer to that of Carr et al. (2023) than that of Voit et al. (2024a). At the high mass end, the energy loading becomes small, and so the SMHM relation is set primarily by the TNG BH model, though of course the early evolution of such objects may also be different.
3.2.3 Stellar mass – black hole mass relation